# Towards Vision Enhancing LLMs: Empowering Multimodal Knowledge Storage and Sharing in LLMs

### v1

@@ 3.2 Multimodal Knowledge Collaboration
After gaining visual information storage inside LLMs, we need consider how to realize multimodal knowledge collaboration during generation. Regarding pretrained MVM and MLP in LLMs as visual and textual experts respectively, we propose a soft mixtures-of-multimodal experts (MoMEs) approach to achieve multimodal knowledge utilization at the token level.
Mixtures-of-Multimodal Experts (MoMEs). To speed up training process, as the bottom part shown in Figure 3, we freeze MVM and other parameters of LLMs, applying Low-Rank Adaption (Hu et al., 2021) (LoRA) to tow-modality experts: MVM and MLP. We denote the inputs tokens for one sequence inputted to MoMEs by $\mathbf{X}\in\mathbb{R}^{m\times d}$, where $m$ is the number of tokens and $d$ is their dimension. The computed process for visual and language knowledge expert could be given in
@@
$LoRA(W_{0}):=W_{0}X+\Delta WX=W_{0}x+BAX,$
where $B,A$ are learnable parameters added for each pretrained weight of visual and textual experts. LoRA-MVM and LoRA-MLP(X) represent original knowledge experts equipped with additional LoRA calculation. By doing so, the training process is efficient because of doing not update the overall parameters of experts.
@@
where $S\in R^{X\times 2}$ and the final dimension is normalized with Softmax calculation. The output of each block in LLMs is denoted to $\mathbf{h}_{o}$.
@@
Table 1: Zero-shot model performances on natural language processing benchmarks. Models with ${}^{\dagger}$ indicate that their SFT or RLHF tuned data are unknown or unused in our work. “INST-LoRA” refers to applying the widely-used LoRA technical to fine-tune LLMs with same text-only instruction data. “Multimodal-SFT” represents the multimodal instruction-following data. “Avg.Score” shows the average evaluation score on the total tasks. Bold and underlined numbers refer to the best and second-best performance for comparative model variants of Llama-2-7b/13b, respectively.
@@
${}^{\dagger}$ (Touvron et al., 2023) 37.02 37.80 49.46 42.89 67.29 27.45 47.69 44.23
${}^{\dagger}$ (Chiang et al., 2023) 58.21 38.82 55.85 55.46 67.01 37.02 50.96 51.90
[…]
@@
w/o Multimodal-SFT & MoMEs 54.81 68.21 62.25 54.0 67.95 35.08 46.50 55.54
${}^{\dagger}$ (Touvron et al., 2023) 31.62 36.83 42.37 35.3 64.90 23.53 37.05 38.82
${}^{\dagger}$ (Chiang et al., 2023) 42.58 67.58 39.71 38.2 55.62 29.32 38.94 44.56
[…]
@@
w/o Multimodal-SFT & MoMEs 42.84 70.46 55.42 37.0 60.71 25.27 37.71 47.06
3.3 Training
In the first stage, the size of used image-text pairs are about 2.3M from CC3M (Changpinyo et al., 2021), COCO Captioning (Chen et al., 2015), and Flick-30k (Plummer et al., 2015). To achieve multimodal knowledge collaboration, as shown in Figure 1, we use text-only and image-text instruction-following data to train the overall architecture. The added modular visual memory and LLMs are frozen during training. We use widely-used instruction data including: high-quality natural language processing tasks from Flan-T5 (Chung et al., 2022), complex instruction-finetuning data from WizardLLMs (Xu et al., 2023), and multimodal instruction data LLaVAR (Zhang et al., 2023), which totally consists of 1.5M text-only and 166k image-text instruction tuning data.
4 Experiments
4.1 Datasets
Natural Language Processing Benchmarks. We use seven text-only downstream datasets to comprehensively evaluate MKS2, which consists of physical world knowledge-relevant datasets and basic ability assessment benchmark MMLU (Hendrycks et al., 2021). We use multiple choice question answering tasks that can benefit from visual knowledge: PIQA (Bisk et al., 2020) that requires physical commonsense reasoning, Commonsense QA (CSQA) (Talmor et al., 2019) for evaluating the commonsense reasoning capability of models, OpenBook QA (OBQA) (Mihaylov et al., 2018) that requires multi-step reasoning, use of additional common and commonsense knowledge, and rich text comprehension, RiddleSense (RS) (Lin et al., 2021) for complex understanding of figurative language and counterfactual reasoning skills, Social IQA (Sap et al., 2019) which focuses on physical or taxonomic knowledge for testing social commonsense intelligence, StrategyQA (Geva et al., 2021) that needs the reasoning steps should be inferred using a strategy.
Table 2: Zero-shot performances on some multimodal datasets.“NumImg” represents the total number of images contained in the pretraining stage. The size of input image is always ${}^{2}$ for the following models. “${}^{{\ddagger}}$” indicates that the corresponding model employs the training samples of following evaluation benchmarks such as VQAv2, OK-VQA, and OCR-VQA, leading to unfair comparison.
@@ 4.1 Datasets
Flamingo (Alayrac et al., 2022) $>$1B 49.2 41.2 19.3 27.8 29.0 5.0
MiniGPT-4 (Vicuna-7b) (Zhu et al., 2023) 5M 44.3 32.1 14.0 11.5 18.7 3.0
LLaVA (Vicuna-7b) (Liu et al., 2023a) 0.6M 53.5 47.4 22.1 11.4 28.9 4.5
LLaVAR (Vicuna-13b) (Zhang et al., 2023) 1M - - 30.2 23.4 39.5 6.2
OFA-Large (Wang et al., 2022a) 20M 40.2 19.3 - - - -
FROMAGe (OPT-6.7b) (Koh et al., 2023) 3.3M 44.1 20.1 - - - -
${}^{{\ddagger}}$ (Ye et al., 2023) 11B - - 29.3 28.6 40.3 6.9
${}^{{\ddagger}}$ (FlanT5XL) (Liu et al., 2023a) 129M 62.6 50.1 23.9 39.7 33.1 3.8
BLIP-2 (OPT-6.7b) (Li et al., 2023a) 129M 50.1 36.4 13.4 10.6 21.2 0.8
BLIP-2 (FlanT5XL) (Li et al., 2023a) 129M 42.8 25.6 15.8 26.6 25.2 2.9
BLIP-2 (FlanT5XXL-11B) (Li et al., 2023a) 129M 45.4 27.8 21.7 30.7 32.2 4.9
MKS2-Llama-2-13b 2.3M 54.4 45.1 23.4 35.7 34.2 6.7
MKS2-Llama-2-7b 2.3M 53.3 42.1 22.3 25.2 33.1 6.7
w/o Text-SFT 2.3M 50.2 40.8 21.5 36.5 34.2 7.4
w/o Text-SFT & MoMEs 2.3M 50.1 41.2 21.4 35.3 34.3 7.3
Image-Text Understanding Benchmarks. To evaluate the multimodal capability of our proposed model, we introduce six classical Visual Question Answering (VQA) datasets: VQAv2 (Antol et al., 2015), OK-VQA (Marino et al., 2019), ST-VQA (Biten et al., 2019), OCR-VQA (Mishra et al., 2019), TextVQA (Singh et al., 2019), and DocVQA (Mathew et al., 2021). VQAv2 is a classic open-world VQA dataset, containing more than 1 million samples. Scene Text Visual Question Answering (STVQA) consists of 31,000+ questions across 23,000+ images collected from various public datasets. The OCRVQA dataset includes more than 1 million question-answer pairs that cover over 207,000 book cover images. The TextVQA dataset consists of over 45,000 questions related to text on more than 28,000 images selected from specific categories of the OpenImages dataset. DocVQA is a comprehensive dataset comprising 12,767 document images with diverse types and content, accompanied by over 50,000 questions and answers. For datasets containing far more than 5000 image-question pairs, we selected the first 5000 pairs for test, similar to Liu et al. (2023c).
4.2 Comparing Models
The comparing models mainly comprise three types of open-source large language models Llama-2: SFT Llama-2, RLHF-tuned Llama-2, and recently proposed MLLMs. To verify the new vision enhancing supervised fine-tuning method MKS2, we present text-only instruction tuned variant Llama-2-7b-INST-LoRA and Vicuna-Llama-2 (Penedo et al., 2023), where we adopt the text instruction data identical to our approach and set $r=16$ for LoRA to train Llama2-7b-INST. Hence, the training parameters of Llama-2-7b-INST-LoRA is similar to the proposed MKS2-Llama-2-7B, about 14M. RLHF-tuned models are are language models that have been trained using a combination of human feedback and reinforcement learning techniques, achieving better performance to understand human instruction and generate high-quality responses. We mainly comparing with Llama-2-7b-chat and Llama-2-13b-chat variants released by Meta. Additionally, to evaluate the multimodal information processing capability of MKS2, we also introduce recently proposed MLLMs as baselines. Flamingo (Alayrac et al., 2022) and OFA (Wang et al., 2022a) are traditionally pretrained visual and language models, which have seen an amount of image-text pairs. BLIP-2 (Li et al., 2023a) is a widely-used visual and language models, achieving remarkable zero-shot performance on downstream image-text understanding tasks. MiniGPT-4 (Zhu et al., 2023), FROMAGe (Koh et al., 2023), mPLUG-Owl (Ye et al., 2023), LLaVR (Liu et al., 2023a) and InstructBLIP (Li et al., 2023a) are multimodal instruction tuned MLLMs, trained with enormous image-text instruction-following data.
4.3 Implementation Details
We take the pretrained Llama-2 version (Touvron et al., 2023) as the backbone of MKS2 and run all models with Adam Optimizer (Kingma & Ba, 2014) on 4 A100-80G GPUs with python environment. All models are trained and tested with Bfloat16 floating-point format. The dimension of the middle layer of the inserted visual memory module is 1/4 of the hidden state size of LLMs. For Llama-2-7b, the total parameters of MVM is about 410 million. During visual information storage, we take the frozen visual encoder and Q-former from BLIP-2-FlanT5-xxl to obtain the image encoding, thus the length of soft image embedding is 32. Additionally, we set the initial learning rate to 1e-4 and train the model for about 2 epochs with warm up steps equaling 5000. The batch size is set to 32 with four-step gradient accumulation for single GPU device. While performing instruction-following learning, we set the batch size, $r$ in LoRA to 3 and 8, respectively, and the max length of input is set to 1024. To tag the position of image embedding, we introduce two learnable tokens $img\text{-}start\right>$ and $img\text{-}end\right>$. Similar to Llama-2-chat, we add [INST] and [/INST] at the starting and ending of inputting text instruction, as like “[INST] Please write a short story about cat and dog [/INST]”. During generation, we set beam sizes to 1 and 4 for text-only and VQA tasks respectively.
4.4 Overall Performances
Performance of vision enhancing LLMs. We present zero-shot model performances on Table 1, aiming to evaluate the instruction-understanding and open-world problem solving abilities of LLMs. We observe that the proposed method MKS2-Llama-2-7B/13B achieves best performances on almost all evaluation datasets, especially on substantially suppressing Llama-2-7b/13b-chat. Compared to powerful Llama-2-7b-INST-LoRA of the same magnitude, MKS2-Llama-2-7b could gains by about 8% on CommensenseQA, 14.5% on OpenBookQA, 16% on PIQA, and 8.6% on RS, respectively. Hence, MKS2 is capable of markedly improving the overall performance on text-only tasks requiring physical world knowledge. Compared to Vicuna-Llama-2 models with all parameters of Llama-2 updating, our approach stands out by requiring being fine-tuned on only a small fraction of parameters ($<$ 0.2% of LLM parameters) while still achieving superior performance on several tasks.
@@ 4.4 Overall Performances
Competitive performances on multimodal benchmarks. We also present the model’s zero-shot performance on the VQA dataset in Table 2. To gain suitable and robust image embeddings, we further fine-tuned the visual mapping network for one epoch and freeze all other parameters, which does not affect any text-only performance of LLMs. We can see that MKS2-Llama-2-7b could achieve comparative performances on open-world and scene text VQA datasets. It’s noteworthy that there was no discernible performance degradation when comparing MKS2 performance on multimodal tasks to that of the original large language model without visual enhancement. This implies that the addition of visual enhancement in LLMs did not lead to a loss in text-related knowledge, while performing LLMs for vision. Moreover, the incorporation of text-only data proves to be a valuable strategy for enhancing the model proficiency in answering open visual questions. While mixed instruction data leads to improvements in open-world question answering, it appears to have a detrimental effect on scene text recognition, possibly due to shifts in training data distribution. Further investigation into the fine-tuning process and data distributions can help optimize MKS2 performance across a wider range of tasks involving both text and images.

### v2

@@ 4.2 Multimodal Knowledge Collaboration
After gaining visual information storage inside LLMs, we need to consider how to realize multimodal knowledge collaboration during generation. Regarding pretrained MVM and MLP in LLMs as visual and textual experts, respectively, we propose a soft mixture-of-multimodal experts approach to achieving multimodal knowledge utilization at the token level.
Mixture-of-multimodal Experts (MoMEs). To speed up the training process, as the bottom part shown in Figure 3, we freeze MVM and other parameters of LLMs, applying Low-Rank Adaption (Hu et al., 2021) (LoRA) to tow-modality experts: MVM and MLP. We denote the tokens of the input for one sequence inputted to MoMEs by $\mathbf{X}\in\mathbb{R}^{m\times d}$, where $m$ is the number of tokens and $d$ is their dimension. As show in Figure 4, the computed process for visual and language knowledge experts could be given in
@@
$LoRA(W_{0})=W_{0}X+\Delta WX=W_{0}x+BAX,$
where $B,A$ are learnable parameters added for each pretrained weight of the visual and textual experts. LoRA-MVM(X) and LoRA-MLP(X) represent original knowledge experts equipped with additional LoRA calculation. By adopting this method, the training process becomes more efficient, as it does not require updating the overall parameters of two experts.
@@
where $S\in R^{X\times 2}$ refers to the weights of each expert, and the final dimension is normalised with Softmax calculation. The output of the MoMEs block in LLMs is denoted to $\mathbf{h}_{M}$.
5 Training and Data Recipes
In the first stage, the size of image-text pairs is about 2.3M (Million) from CC3M Changpinyo et al. (2021), COCO Captioning Chen et al. (2015), and Flick-30k Plummer et al. (2015). To achieve multimodal knowledge collaboration, as shown in Figure 1, we use text-only and image-text instruction-following data to train the overall architecture. The modular visual memory and LLMs are frozen during training. We use widely-used instruction data including natural language processing tasks filtered from Flan-T5 Chung et al. (2022), complex instruction-finetuning data from WizardLLMs Xu et al. (2024), and multimodal instruction data LLaVAR Zhang et al. (2023b), which consists of 1.5M textual and 166k image-text question-answering pairs.
During text-only inference, the model functions as a powerful LLM, with the added capability of recalling knowledge stored in its MoMEs from past multimodal learning. The transition between multimodal and text-only inputs is seamless because the visual processing pathway is an additive, optional module that leaves the core language model intact. In this mode, the MoMEs remain dormant yet accessible, having been integrated into the LLM’s architecture and imbued with knowledge from visual experiences. When faced with a question that requires visual knowledge, e.g., “What is the Eiffel Tower made of ?”, the model can activate the relevant experts within the MoMEs to retrieve the necessary information, analogous to how it activates neurons for purely textual facts. For multimodal inputs, our model functions similarly to traditional MLLMs, utilizing a vision encoder to process the input image.
6 Experiments
6.1 Datasets
6.1.1 Natural Language Processing Benchmarks
We use seven text-only downstream datasets to comprehensively evaluate MKS2, which consists of physical world knowledge-relevant datasets and basic ability assessment benchmark MMLU (Hendrycks et al., 2021). We use multiple choice question answering tasks that can benefit from visual knowledge: PIQA (Bisk et al., 2020) that requires physical commonsense reasoning, Commonsense QA (CSQA) (Talmor et al., 2019) for evaluating the commonsense reasoning capability of models, OpenBook QA (OBQA) (Mihaylov et al., 2018) that requires multi-step reasoning, use of additional common and commonsense knowledge, and rich text comprehension, RiddleSense (RS) (Lin et al., 2021) for complex understanding of figurative language and counterfactual reasoning skills, Social IQA (Sap et al., 2019) which focuses on physical or taxonomic knowledge for testing social commonsense intelligence, StrategyQA (Geva et al., 2021) that needs the reasoning steps should be inferred using a strategy.
6.1.2 Image-Text Understanding Benchmarks
To evaluate the multimodal capability of our proposed model, we introduce six classical Visual Question Answering (VQA) datasets: VQAv2 (Antol et al., 2015), OK-VQA (Marino et al., 2019), ST-VQA (Biten et al., 2019), OCR-VQA (Mishra et al., 2019), TextVQA (Singh et al., 2019), and DocVQA (Mathew et al., 2021). VQAv2 is a classic open-world VQA dataset, containing more than 1 million samples. Scene Text Visual Question Answering (STVQA) consists of 31,000+ questions across 23,000+ images collected from various public datasets. The OCRVQA dataset includes more than 1 million question-answer pairs that cover over 207,000 book cover images. The TextVQA dataset consists of over 45,000 questions related to text on more than 28,000 images selected from specific categories of the OpenImages dataset. DocVQA is a comprehensive dataset comprising 12,767 document images with diverse types and content, accompanied by over 50,000 questions and answers. For datasets containing far more than 5000 image-question pairs, we selected the first 5000 pairs for the test, similar to (Liu et al., 2023c).
6.2 Comparing Models
The comparing models mainly comprise three types of open-source Llama-2 variants: SFT Llama-2, RLHF-tuned Llama-2, and recently proposed MLLMs. To verify the vision-enhancing method MKS2-Llama, we present text-only instruction-tuned variant Llama-2-7b-INST-LoRA and Vicuna-Llama-2 (Penedo et al., 2023), where we adopt the textual instruction data identical to our approach and set $r=16$ for LoRA to train Llama2-7b-INST. Hence, the training parameters of Llama-2-7b-INST-LoRA are similar to the proposed MKS2-Llama-2-7B, about 14M. RLHF-tuned models are language models that are trained using a combination of human feedback and reinforcement learning techniques, achieving better performance in understanding human instructions and generating high-quality responses. Additionally, to evaluate the multimodal information processing capability of MKS2-Llama, we also introduce recent MLLMs as baselines. Flamingo (Alayrac et al., 2022) and OFA (Wang et al., 2022b) are traditionally pretrained visual-language models, which have seen an amount of image-text pairs. BLIP-2 (Li et al., 2023b) is a widely-used visual and language model, achieving remarkable zero-shot performance on downstream image-text understanding tasks. MiniGPT-4 (Zhu et al., 2024), KOSMOS-2 (Peng et al., 2023), ImageBind-LLM (Han et al., 2023), FROMAGe (Koh et al., 2023), mPLUG-Owl (Ye et al., 2023), LLaVR (Zhang et al., 2023b) and InstructBLIP (Dai et al., 2024) are multimodal instruction-tuned MLLMs, trained with enormous image-text instruction-following data. In addition, for the knowledge-based VQA tasks, we introduced two task-specific model: ViperGPT (Surís et al., 2023) with code-based visual reasoning and REVEAL (Hu et al., 2023) with multimodal knowledge retrieval generation.
6.3 Implementation Details
We take the pretrained Llama-2 version (Touvron et al., 2023) as the backbone of MKS2 and run all models with Adam Optimizer (Kingma & Ba, 2014) on 4 A100-80G GPUs with a Python environment. All models are trained and tested with the Bfloat16 floating-point format. The dimension of the middle layer of the inserted visual memory module is 1/4 of the hidden state size of LLMs. For Llama-2-7b, the total parameters of MVM are about 410 million. During visual information storage, we take the frozen visual encoder and Q-former from BLIP-2-FlanT5-xxl to obtain the image encodings, so the length of soft image embedding is 32. Additionally, we set the initial learning rate to 1e-4 and trained the model for about 2 epochs with warm-up steps equaling 5000. The batch size is set to 32 with four-step gradient accumulation for a single GPU device. While performing instruction-following learning, we set the batch size, $r$ in LoRA to 3 and 8, respectively, and the max length of the input is set to 1024. To tag the position of image embedding, we introduce two learnable tokens $\left<img\text{-}start\right$> and $\left<img\text{-}end\right$>. Similar to Llama-2-chat, we add [INST] and [/INST] at the start and end of text instruction, as like “[INST] Please write a short story about cat and dog [/INST]”. During generation, we set beam sizes to 1 and 4 for text-only and VQA tasks, respectively.
6.4 Overall Performance
Performance of vision-enhancing LLMs. We present zero-shot model performances in Table 1, aiming to evaluate the instruction-understanding and open-world problem-solving abilities of LLMs. We observe that the proposed method MKS2-Llama-2-7B/13B achieves the best performance on almost all evaluation datasets, especially on substantially suppressing Llama-2-7b/13b-chat. Compared to powerful Llama-2-7b-INST-LoRA of the same magnitude, MKS2-Llama-2-7b could gain by about 8% on CommensenseQA, 14.5% on OpenBookQA, 14.5% on PIQA, and 6.2% on RS, respectively. Hence, MKS2 is capable of markedly improving the overall performance on text-only tasks requiring physical world knowledge. Compared to Vicuna-Llama-2 models with all parameters of Llama-2 updating, our approach stands out by requiring being fine-tuned on only a small fraction of parameters (0.25% of LLM parameters) while still achieving superior performance on several tasks.
@@ 6.4 Overall Performance
Table 1: Zero-shot model performances on natural language processing benchmarks. Models with indicate that their SFT or RLHF-tuned data are unknown or unused in our work. “INST-LoRA” refers to applying the widely-used LoRA technical to fine-tune LLMs with the same text-only instruction data. “Multimodal-SFT” represents the multimodal instruction-following data. “Avg” refers to the average evaluation score on the total tasks. Bold and underlined numbers refer to the best and second-best performance for comparative model variants of Llama-2-7b/13b, respectively.
@@
KOSMOS-2 (Peng et al., 2023) - - - - 72.9 - - -
Llama-2-13b-chat (Touvron et al., 2023) 37.02 37.80 49.46 42.89 67.29 27.45 47.69 44.23
Vicuna-Llama-2-13b (Chiang et al., 2023) 58.21 38.82 55.85 55.46 67.01 37.02 50.96 51.90
Llama-2-13b-INST-LoRAr=16 57.68 63.73 63.80 58.6 71.98 38.51 46.70 57.28
@@
w/o (Multimodal-SFT & MoMEs) 54.81 68.21 62.25 54.0 67.95 35.08 46.50 55.54
Llama-2-7b-chat (Touvron et al., 2023) 31.62 36.83 42.37 35.3 64.90 23.53 37.05 38.82
Vicuna-Llama-2-7b (Chiang et al., 2023) 42.58 67.58 39.71 38.2 55.62 29.32 38.94 44.56
Llama-2-7b-INST-LoRAr=16 41.93 74.10 54.65 39.4 53.42 27.01 38.68 47.02
@@
w/o (Multimodal-SFT & MoMEs) 42.84 70.46 55.42 37.0 60.71 25.27 37.71 47.06
Table 2: Model performances on multimodal datasets.“NumImg” represents the total number of images contained in the pretraining stage. The size of the input image is always 2242 for the following models, and answers are directly generated, except for LLaVA-v1.5 with the image size of 3362. “” indicates that the corresponding model employs the training sets of the following evaluation benchmarks such as VQAv2, OK-VQA, and OCR-VQA. MKS2-Llama-2-7b and LLaVA-v1.5 employ the same size of visual instruction tuning data, which is used to compare the performance of introducing the mixture-of-expert architecture.
@@
Flamingo (Alayrac et al., 2022) >>1B 49.2 41.2 19.3 27.8 29.0 5.0 28.6
MiniGPT-4 (Vicuna-7b) (Zhu et al., 2024) 5M 44.3 32.1 14.0 11.5 18.7 3.0 20.6
OFA-Large (Wang et al., 2022b) 20M 40.2 19.3 - - - - -
FROMAGe (OPT-6.7b) (Koh et al., 2023) 3.3M 44.1 20.1 - - - - -
mPLUG-Owl (Ye et al., 2023) 11B - - 29.3 28.6 40.3 6.9 -
KOSMOS-2 (Peng et al., 2023) >>2B 45.6 - - - - - -
ViperGPT (>>11B) (Surís et al., 2023) >>129M - 48.1 - - - - -
ImageBind-LLM (Chinese-LLama-7B) (Han et al., 2023) >>3M - 51.7 15.5 23.2 24.0 4.0 -
REVEAL (2B) (Hu et al., 2023) >>12M 59.1 - - - - -
InstructBLIP (FlanT5XL) (Dai et al., 2024) 129M 62.6 50.1 23.9 39.7 33.1 3.8 46.0
BLIP-2 (OPT-6.7b) (Li et al., 2023b) 129M 50.1 36.4 13.4 10.6 21.2 0.8 22.1
BLIP-2 (FlanT5XL) (Li et al., 2023b) 129M 42.8 25.6 15.8 26.6 25.2 2.9 23.2
BLIP-2 (FlanT5XXL-11B) (Li et al., 2023b) 129M 45.4 27.8 21.7 30.7 32.2 4.9 27.1
LLaVAR (Vicuna-13b) (Zhang et al., 2023b) 1M 54.2 44.9 30.2 23.4 39.5 6.2 33.1
MKS2-Llama-2-13b 2.3M 54.4 45.1 28.4 35.8 37.2 6.8 34.6
[1pt/1pt] LLaVA (Vicuna-7b) (Liu et al., 2023a) 0.6M 53.5 43.2 22.1 11.4 28.9 4.5 27.3
LLaVAR (Vicuna-7b) (Zhang et al., 2023b) 1M 51.3 40.6 28.9 24.9 35.8 6.2 31.3
MKS2-Llama-2-7b 2.3M 53.3 42.1 22.3 25.2 33.1 6.7 30.5
w/o Text-SFT 2.3M 50.2 40.8 21.5 36.5 34.2 7.4 31.7
w/o (Text-SFT & MoMEs) 2.3M 50.1 41.2 21.4 35.3 34.3 7.3 31.6
[1pt/1pt] LLaVA-v1.5 (Vicuna-7b) (Liu et al., 2024a) 1M 78.4 59.8 51.3 53.1 53.1 22.4 53.0
MKS2-Llama-2-7b 1M 83.3 64.7 51.6 54.8 53.2 22.8 55.1
Figure 5: An illustration of cases generated by the pretrained MVM, where we add the generation loss as the supervision object. It’s important to highlight that we assess image generation quality primarily to verify whether the added MVM effectively stores text-visual knowledge and establishes the connection between language and vision. Fortunately, the MVM can connect the language to their corresponding imagination.
Table 3: Top@1 accuracy of text-to-image retrieval on Image Retrieval (Flick 30k) and Image captioning Quality. “” indicates that the corresponding model is trained on the corresponding training set.
[…]
CLIP (Hafner et al., 2021) 38.2 68.7
ALIGN (Li et al., 2021) - 75.7
FILIP (Yao et al., 2021) - 75.0
E2E-VLP (Xu et al., 2021) 36.2 73.6
Florence (Yuan et al., 2021) - 76.7
Oscar (Li et al., 2020) 41.7 57.5
CoCa (Yu et al., 2022) 40.9 76.8
OFA (Wang et al., 2022b) (20M Images) 44.9 82.2
mPLUG (Ye et al., 2023) 46.5 88.4
GIT (Wang et al., 2022a) (0.8B Images) 44.1 -
BLIP-2(Li et al., 2023b) (129M Images) 42.4 89.7
Our Model (2.4M Images) 40.0 80.0
@@
Competitive performance on multimodal benchmarks. We also present the model’s zero-shot performance on the VQA dataset in Table 2. To gain suitable and robust image embeddings, we further fine-tuned the visual mapping network for one epoch and froze all other parameters, which does not affect any text-only performance of LLMs. Our model demonstrates competitive performance across diverse VQA benchmarks. Comparing MKS2 w/o Text-SFT to the base MLLM (w/o (Text-SFT & MoMEs) ) reveals no discernible performance degradation on multimodal tasks, indicating that the visual enhancement successfully preserves the model’s original textual knowledge. Furthermore, MKS2 outperforms several prior specialized approaches-including those focused on visual feature injection (e.g., ImageBind-LLM), code-based reasoning (e.g., ViperGPT), and multimodal retrieval (e.g., REVEAL). This superior performance validates our novel methodology of using a language-centred learning strategy and a visual memory module to internalize multimodal knowledge directly within the LLM’s parameters. This implies that the addition of visual enhancement in LLMs did not lead to a loss in text-related knowledge while performing LLMs for vision. Moreover, the incorporation of text-only data proves to be a valuable strategy for enhancing the model’s proficiency in answering open visual questions. While beneficial for QA, mixed instruction data harms scene text recognition, likely due to training data shifts. Optimizing MKS2’s performance requires investigating fine-tuning and data distributions.
Table 4: Ablation Experiments on five datasets requiring common sense. We explore various data and visual memory sizes to check model performances. All models are built upon Llama-2-7b, where MKS2 signifies MKS2 w/o Multimodal-SFT. "-AM-BM" indicates that it includes an "A" size of visual memory and was trained using "B" number of image-text pairs.
[…]
Llama-2-7b-chat 31.62 35.3 64.90 23.53 42.37
Vicuna-Llama-2-7b 42.58 38.2 53.42 29.32 39.71
MKS2-410M-2.3M 44.06 50.5 67.10 28.99 57.72
[…]
MKS2-810M-2.3M 43.78 48.6 66.97 28.89 57.21
MKS2-810M-12.3M 46.56 49.8 68.10 30.65 57.84
6.5 Visual Storage Analysis
This section mainly conducts further studies to analyze the effectiveness of modular visual memory (MVM). we first validate the effectiveness of the MVM by comparing the performance of two configurations: MKS2 w/o (Multimodal-SFT & MoMEs) and MKS2 w/o Multimodal-SFT, where “w/o (Multimodal-SFT & MoMEs) ” shows that we do not utilize the pretrained MVM module. When the MVM is employed, the significant improvement in performance attests to its capacity for external information storage. Given that the MVM is exposed solely to visual information during pretraining, we deduce that it stores visual information effectively.
In Table 3, we can see that our model trained with 2.3M achieves certain image description generation and image retrieval capabilities. This indicates that using a large language model as a text understanding encoder could attain better image-text connection capabilities. This might be because the large language model has larger parameters and stronger language understanding capabilities. However, compared to multimodal encoders pre-trained with tens of millions or hundreds of millions of image-text pairs and supervised fine-tuned multimodal models, there is a large performance gap. Nevertheless, increasing the pre-training data of image-text pairs can make up for this performance gap. Although this is not the focus of this paper, our main objective is to enhance the knowledge scope of the large language model and verify that the proposed MVM is capable of storing visual information.
Additionally, we demonstrated the ability of our model to generate images based on text prompts after the first stage of pre-training in Figure 5. From the cases displayed in the figure, we observed that when image generation is added as the loss of the first stage, our model could still adapt and generate relatively good images. In other words, it can imagine relevant visual content based on the text information. The experimental results in Table 3 and the image generation quality in Figure 5 indicate that the introduced language-centric learning method can inspire MVM to store visual information and establish the connection between images and text. By adding a simple MLP to each layer of the pre-trained large language model, we could expand the knowledge storage surface of the large language model, and the relevant parts could be activated through a language-centric learning strategy.
To further examine the MVM’s capability for open-world visual information storage, we varied the number of images in our experiments to assess whether the MVM could accommodate different scales of visual information. This variation directly influences the model’s downstream performance. The experimental outcomes, detailed in Table 4, indicate that the downstream performance fluctuates with the size of the training image dataset, under consistent experimental setups. This variation underscores the MVM’s adaptability to different scales of visual information.